OpenStudy (anonymous):

A monument is built out of toothpicks. There are 25 rows and each row has eight more toothpicks than the row above. The top row has five toothpicks.

OpenStudy (anonymous):

ok whats the question

OpenStudy (anonymous):

Part 1: How many toothpicks are on the bottom row?

OpenStudy (anonymous):

Part 2: Using complete sentences, explain the procedure taken to answer this question. (4 points)

OpenStudy (anonymous):

@prowrestler hey are you there?

OpenStudy (anonymous):

is this on a homework

OpenStudy (anonymous):

Yeah I just need help on this question though- every time I do I get a new answer

OpenStudy (anonymous):

This will form an AP The first row has 5 Next let last one has L toothpicks... Now formula for last term in AP is... L= a + (n-1)d Lastly.....d=8 a=5 n=25 L= 5 + 24*8 =197

OpenStudy (anonymous):

Thank you!